OpenStudy (anonymous):

Point slope form with these points: (10, 5) and (12, 12)

OpenStudy (anonymous):

I need an equation ;)

OpenStudy (anonymous):

i think the slope of the line between the opoints is 7/2, as it is \[ \frac{change in y}{change in x} \]

OpenStudy (anonymous):

so then you do y-5 = (7/2)(x-10) 2y-10=7x-70 2y-7x+60=0

OpenStudy (anonymous):

yeah \(\frac{12-5}{12-10}\)

OpenStudy (anonymous):

So, the equation is: 2y - 7x + 60 = 0

OpenStudy (anonymous):

equation is usually expressed in this form y=kx+b

OpenStudy (anonymous):

yes, or you can write it as \[y = \frac{7}{2}*x - 60\]

OpenStudy (anonymous):

that is the equation in the form tomas said, either one is ok

OpenStudy (anonymous):

you forgot to divide 60 by 2

OpenStudy (anonymous):

\[y = \frac{7}{2}*x - 30\]

OpenStudy (anonymous):

Okay. Thanks Tomas.A!!! :)

OpenStudy (sasogeek):

point slope? is that the point slope form.... i thought that was intercept slope form... i'm lost

OpenStudy (anonymous):

i thinks it's same but im not sure since i didn't learn maths in english

OpenStudy (sasogeek):

point slope => y2-y1=m (x2-x1)

OpenStudy (anonymous):

so she is asked to find point slope which is m and then she also asked for equation?

OpenStudy (sasogeek):

(10, 5) and (12, 12) m=(12-5)/(12-10) m=7/2 point slope form of equation of line on which the 2 points lie (y-5)=(7/2) * (x-10)

OpenStudy (sasogeek):

(y-5)=(7/2) * (x-10) y-5=(7/2)x - (7/2 * 10) y-5=(7/2)x - 7*5 y=7/2x - 30

OpenStudy (anonymous):

yeah 08bkrishna wrote like that

OpenStudy (sasogeek):

point slope form and intercept slope form are quite confusing, i normally go by y=mx+c but i only recently learned the difference in the 2 equation forms

OpenStudy (anonymous):

So, what is the equation! :)

OpenStudy (sasogeek):

point slope form of equation of line on which the 2 points lie (y-5)=(7/2) * (x-10) if you keep simplifying it down, you'll result with the intercept slope form.

OpenStudy (anonymous):

that's the final answer \[y = \frac{7}{2}x - 30\]

OpenStudy (sasogeek):

the final answer is an intercept slope form... i believe the question was clearly asking for point slope, not intercept slope http://www.purplemath.com/modules/strtlneq2.htm

OpenStudy (sasogeek):

the point slope from isn't a simplified equation, if you go ahead to simplify it, you'll end up with the intercept slope form, that's the confusing part as to whether you should keep solving it but once u have the point slope form and that's what is being required of, stop there